OpenStudy (anonymous):

An arrow shot horizontally from a cliff at 15 m/s lands 30 m away. How how high is the cliff ?

OpenStudy (anonymous):

Projectile motion in 2D (x,y plane) : It is fairly lengthy to derive these equations:

OpenStudy (anonymous):

I cannot make an equation now...

OpenStudy (anonymous):

x=( v cos alpha)t y=(v sin alpha)t -1/2 g t^2 Where alpha is the angle you shoot the arrow. v is the initial velocity t is time g is the universal gravitational constant

OpenStudy (anonymous):

does it have to be so complex?

OpenStudy (anonymous):

The answer for this question is 19.6cm height i dont know how to solve .

OpenStudy (anonymous):

you need the right formula i think

OpenStudy (anonymous):

There will be so many simplification as you are shooting the arrow horizontally. Cos alpha=1 Sin alpha =0

OpenStudy (anonymous):

x=vt y=-1/2 g t^2

OpenStudy (anonymous):

30=15 t Thus t=2

OpenStudy (anonymous):

y=-1/2 g 2^2=-2g=-2*9.80=-19.6

OpenStudy (anonymous):

I am correct ?

OpenStudy (anonymous):

Well you just had the solution :) And yes that is it

OpenStudy (anonymous):

well done andras

OpenStudy (anonymous):

Thanks for the help :)

OpenStudy (anonymous):

Somehow I could not create proper equations, so it does not look so neat

OpenStudy (anonymous):

Andras I am now your FAN :)

OpenStudy (anonymous):

Yeah I lvld up :) 70 now

OpenStudy (anonymous):

Hahaha :)

OpenStudy (anonymous):

Andras the solution is correct ?